An curve has equation y =12/3 -2x

Find dy/dx


1
Expert's answer
2021-07-29T09:48:32-0400

"y = \dfrac{12}{3-2x}"


"u = 12
\\
du/dx= 0
\\
v = 3- 2x
\\
dv /dx = -2"


from, "\dfrac{dy}{dx}=\dfrac{V\dfrac{du}{dx}- U\dfrac{dv}{dx} }{V²}"


"\dfrac{dy}{dx} =\dfrac{(3-2x)(0)- 12(-2) }{(3-2x)²}"


"\dfrac{dy}{dx} =\dfrac{24}{(3-2x)²}"